There is a pressure because of the ocean water at a 20 m distance underneath the surface, what is the pressure? Take 1.035 g/cm3 density for the ocean water and 1 atm = 101.33 kPa.

P1 = Unit in kPa

There is a pressure because of the ocean water at a depth of 200 m, what is the pressure?

P2 = Unit in Kpa

At what distance underneath the surface the pressure because of the water is 140 atm?

d = Unit in m
Answer #1

(1) Pressure = Depth × Density× g

= 20× 1035× 9.8 = 202860 Pa

= 202.86 kPa

(2) similarly, at 200 m

Pressure = 200× 1035× 9.8=2028600 Pa

= 2028.6 kPa

(3) Depth = Pressure/ (g× density)

= (140×101330)/(9.8× 1035) =1398.62 m